OpenStudy (anonymous):

solve the initial value problem, xy'=y+x^3, y(2)=5.

OpenStudy (anonymous):

xy'=y+x^3 xdy/dx -y =x^3 xdy-ydx =x^3 dx (xdy-ydx)/X^2 =xdx d(y/x) = xdx... now integrate both sides...

OpenStudy (anonymous):

integrating we have y/x = x^2/2 +c

OpenStudy (anonymous):

what about y(2)=5?

OpenStudy (anonymous):

and how does xdy-ydx become d(y/x) instead of d(x/y)? and why does x^2 disappear?

OpenStudy (anonymous):

you need to find c

OpenStudy (anonymous):

put x=2 and y=5 to find c